Hi can someone please help me.

If 6 is subtracted from the third of three consecutive odd integers and the result is multiplied by 2​, the answer is 23 less than the sum of the first and twice the second of the integers. Find the integers.

Answer by rothauserc(4718)

let x be the first odd integer, then we have
:
x, x+2, x+4
:
(x+4 - 6) * 2 = x + 2(x+2) - 23
:
2x - 4 = 3x - 19
:
x = 15
:
***************************
the integers are 15, 17, 19
***************************
: